Determine the concentration of an HBr solution if a 55.00 mL aliquot of the solution yields 0.6643 g AgBr when added to a solution with an excess of Ag ions. The Ksp of AgBr is 5.0
Answer #1

Mol wt of AgBr = 187.7

moles of AG = 0.6643/187.7 = = 0.00354

=> moles of Br = = 0.00354

and molarity = = 0.00354/0.055 = 0.06436 M